MICROWAVE SPECTROSCOPY OF METAL MONOCYANIDE, MCN (M=Ag,Au)

\underline{EMI Y. OKABAYASHI},, Department of Chemistry, Faculty of Science, Shizuoka University, Oya 836, Shizuoka, Japan; TOSHIAKI OKABAYASHI, Center for Instrumental Analysis, Shizuoka University; FUMI KOTO, and MITSUTOSHI TANIMOTO, Department of Chemistry, Faculty of Science, Shizuoka University, Oya 836, Shizuoka, Japan.

The rotational spectrum of MCN (M=Ag, Au) was observed by employing a source-modulation mi crowave spectrometer. The MCN species was generated in a dc glow discharge through the mixture of CH_{3}CN and Ar by a sputtering reaction with an metal target. The spectrum of MCN was observed in the ground and ν_{2} vibrational excited states. The molecular constants derived from the observed transition frequencies were compared with values of theoretical calculations reported in literature.